how many mL of H2SO4 are neutralized if it takes 28.03 mL of 3.55 M NaOH to neutralize a 0.25 sample of H2SO4



1
Expert's answer
2021-05-26T07:08:23-0400

H2SO4(aq)+2NaOH(aq)"\to" Na2SO4(aq)+2H2O(l)

Moles of NaOH=(28.03*3.55)/1000=0.0995065moles

If 2=0.0995065moles

"\therefore" 1=(1/2)*0.0995065

=0.0497533moles

Volume of H2SO4=(0.0497533moles*1000)/0.25

=199.013mL